### v1

# Interacting chiral fermions on the lattice with matrix product operator norms

Abstract
We develop a formalism for simulating one-dimensional interacting chiral fermions on the lattice without breaking any local symmetries by defining a Fock space endowed with a semi-definite norm defined in terms of matrix product operators. This formalism can be understood as a second-quantized form of Stacey fermions, hence providing a possible solution for the fermion doubling problem and circumventing the Nielsen-Ninomiya theorem. We prove that the emerging theory is hermitian by virtue of the fact that it gives rise to a hermitian generalized eigenvalue problem and that it has local features as it can be simulated using tensor network methods similar to the ones used for simulating local quantum Hamiltonians. We also show that the scaling limit of the free model recovers the chiral fermion field. As a proof of principle, we consider a single Weyl fermion on a periodic ring with Hubbard-type nearest-neighbor interactions and construct a variational generalized DMRG code demonstrating that the ground states of the system for large system sizes can be determined efficiently.
I Introduction and main results
Simulating chiral quantum field theories on the lattice is a central problem in both high energy and condensed matter physics. The Nielsen-Ninomiya theorem [1, 2] prohibits the existence of a local symmetry-preserving lattice model with the correct continuum limit due to the appearance of unphysical low-energy modes (“doublers”), which couple to the physical ones in the presence of interactions. Any local lattice formulation such as Wilson fermions [3], Ginsparg–Wilson fermions [4] or staggered fermions [5] either violate an internal or lattice symmetry. The only lattice models that exhibit chiral fermions without breaking any other symmetries are non-local theories [6, 7], which is unsatisfactory. The problem can be mitigated by defining the theory to live on the edge of a higher-dimensional theory [8, 9, 10], but this leads to a much larger simulation cost in practice, and some symmetries are still broken.
In this Letter, we provide an alternative solution that preserves all symmetries without the need to increase the dimensionality of spacetime by endowing our Hilbert space with a semi-definite metric in the form of a matrix product operator (MPO) [11, 12]. Intuitively, this metric mimics the role of the higher-dimensional bulk theory, but it is constructed such that locality is preserved. For the sake of clarity, we will limit our discussion to the one-dimensional case, but all arguments extend to higher dimensions.
Our work is based on the ground-breaking paper of Stacey [13], which solved the fermion doubling problem in the single-particle case by making use of the implicit (versus the usual explicit) way of discretizing the differentials in the field equations. We can also reinterpret this construction from a finite-elements perspective, as discussed below. Indeed, following [14, 15], Stacey’s theory amounts to working with non-orthogonal orbitals and leads to a mapping of the stationary solutions of a single-particle Weyl equation
$i\tfrac{\partial\ }{\partial t}\psi(x,t)=-i\tfrac{\partial\ }{\partial x}\psi(x,t)$ (1)
to a generalized eigenvalue problem of $L\times L$ tridiagonal matrices
$H|\phi\rangle=\lambda N|\phi\rangle$ (2)
$H=\begin{pmatrix}\!0&\!\!\tfrac{i}{2}&\!\!0&\cdots&\!\!-\tfrac{i}{2}\\ \!-\tfrac{i}{2}&\!\!0&\!\!\ddots&\ddots&\!\!\cdots\\ \!0&\!\!\ddots&\!\!\ddots&\ddots&\!\!0\\ \!\vdots&\!\!\ddots&\!\!\ddots&\ddots&\!\!\tfrac{i}{2}\\ \!\tfrac{i}{2}&\!\!\cdots&\!\!0&-\tfrac{i}{2}&\!\!0\end{pmatrix}\!\!,\,\,N=\begin{pmatrix}\tfrac{1}{2}&\tfrac{1}{4}&\!\!0&\cdots&1\\ \tfrac{1}{4}&\tfrac{1}{2}&\!\!\ddots&\ddots&\vdots\\ 0&\ddots&\!\!\ddots&\ddots&0\\ \vdots&\ddots&\!\!\ddots&\ddots&\tfrac{1}{4}\\ \tfrac{1}{4}&\cdots&\!\!0&\tfrac{1}{4}&\tfrac{1}{2}\end{pmatrix}$
where $N$ encodes the overlap between the different orbitals. As both matrices are circulant, they are diagonal in the Fourier basis. The generalized eigenvalues can hence be labeled by the possible momentum $k\!=\!\tfrac{2\pi n}{L}$ and are the ratios of those diagonal entries:
$\lambda_{k}=\tfrac{2\sin(k)}{1+\cos(k)}=2\tan(\tfrac{1}{2}k).$ (3)
The essential trait of this solution is that the unwanted zero-modes of $H$ at $k=pmpi$ are turned into poles by virtue of the norm of the associated wavefunction being singular: the metric $N$ has a zero mode, which breaks the continuity of the dispersion relation (which lies at the origin of fermion doubling).
In this paper, we extend Stacey’s formalism to the many-body case, or in other words, we build a second-quantized version of it. The Fock space is spanned by (unnormalized) vectors
$|n_{1},n_{2},\cdots,n_{L}\rangle=(a_{1}^{\dagger})^{n_{1}}(a_{2}^{\dagger})^{n_{2}}...(a_{L}^{\dagger})^{n_{L}}|\Omega\rangle$
with $n_{i}\in\{0,1\}$, $a_{i}|\Omega\rangle=0$ and commutation relations
$\{a_{i},a_{j}^{\dagger}\}$ $=N_{ij},$$\{a_{i},a_{j}\}=\{a_{i}^{\dagger},a_{j}^{\dagger}\}$ $=0$ (4)
with $N$ the norm-matrix of the single-particle case: this choice indeed reproduces Stacey’s results in the single-particle sector. A crucial feature of the usefulness of this Fock space representation is contained in the following fact: the overlaps between different Fock states are given by a matrix product operator $\tilde{N}$ of bond dimension $\chi=4$. This implies that we can easily calculate norms (as well as expectation values, as will turn out) of matrix product state wavefunctions of the form
$|\psi\rangle=\sum_{n_{1}n_{2}\cdots}\underbrace{\mathrm{Tr\,}\left(A_{1}^{n_{1}}A_{2}^{n_{2}}...A_{L}^{n_{L}}\right)}_{|\phi\rangle}|n_{1},n_{2},...,n_{L}\rangle$
by contracting the tensor network $\langle\phi|\tilde{N}|\phi\rangle$. The physical meaning of this semidefinite norm becomes quite transparent by imagining that the chiral theory lives on the edge of a topological model in 2+1 dimensions, and more specifically of a gapped bulk theory that is represented by a projected entangled-pair state (PEPS) [16, 12]. The edge modes can then be represented by the dangling indices of the PEPS on the border [17], and the norm of such states is determined by the entanglement spectrum of the bulk theory (see figure 1). In [18, 19], it was demonstrated that entanglement Hamiltonians of chiral PEPS also exhibit zero-modes similar to the ones represented here. Nevertheless, we do not need to take such a higher-dimensional theory into account in our simulations, which differentiates our approach from the one in [9, 10].
Figure 1: The matrix product operator fixed point of a double layer PEPS (green) induces a matrix product operator norm (white) for matrix product states (blue) living on the virtual degrees of freedom of the PEPS.
The Hamiltonian, just as any other observable, is expressed in terms of conjugate creation/annihilation operators $b_{i}$. We will simulate the following Hamiltonian:
$\hat{H}=J\sum_{n}\left(i\hskip 2.84526ptb_{n}^{\dagger}b_{n+1}+c.c\right)+U\sum_{n}b_{n}^{\dagger}b_{n+1}^{\dagger}b_{n+1}b_{n},$ (5)
The $b_{i}$ also annihilate the vacuum $|\Omega\rangle$ and satisfy the canonical commutation relations
$\{b_{i},a_{j}^{\dagger}\}$ $=\delta_{i,j}.$
The knowledge of this (anti)commutation rule is sufficient to calculate expectation values of local observables efficiently using the tensor network formalism. Indeed, the Hamiltonian $\hat{H}$ can similarly be retracted onto |$\ket{\phi}$, which we denote as $\tilde{H}$. This retraction furthermore preserves hermiticity so that $\tilde{N}$ and $\tilde{H}$ are Hermitian matrices that give rise to a Hermitian generalized eigenvalue problem with a real spectrum. In the single-particle sector, it reduces to Stacey’s formulation in Eq. (2). But we can now also target the interacting many-body problem, as illustrated in Fig. 2, showing the full generalized spectrum of the above Hamiltonian on a lattice with 11 sites, together with the associated momenta. In the case of non-interacting fermions ($U=0$), the ground state is simply obtained by filling the Fermi sea, which yields a ground state energy equalling the sum of the negative single-particle generalized eigenvalues.
Figure 2: The full generalized spectrum of a chiral fermion in the presence of a Hubbard-type interaction (5) on 11 sites with $n$ particles with $J=U=1$. In the single-particle sector, we recover the correct single-particle spectrum.
As derived in the next section, the dynamics of the system is governed by a generalized Schrödinger equation
$i\tilde{N}\tfrac{\partial\ }{\partial t}|\phi(t)\rangle=\tilde{H}|\phi(t)\rangle$ (6)
Under these dynamics, both $\langle\phi(t)|\tilde{N}|\phi(t)\rangle$ and $\langle\phi(t)|\tilde{H}|\phi(t)\rangle$ are constants of motion. Since $\tilde{H}$ and $\tilde{N}$ are both $U(1)$ invariant, we can use the gauging map for quantum states [20, 21, 22] to gauge the $U(1)$ symmetry of this model. Given that we are dealing with a single right-moving Weyl fermion, however, this $U(1)$ symmetry possesses an anomaly in the quantum field theory, which prohibits its promotion to a gauge symmetry. This apparent contradiction might be resolved by considering that these generalized eigenvalue problems naturally live on the boundary of a higher dimensional bulk theory, as shown in Fig. 1. This is similar to the constructions in [9, 10], where the norm MPO $\tilde{N}$ in our generalized eigenvalue problem plays the role of the bulk theory.
II Action principle and quantization
To motivate the second quantization prescription and the resulting generalized Schrödinger equation, we start from the action principle that gives rise to the Weyl equation (1). Doing so will also enable us to provide another interpretation for the generalized eigenvalue problem coming out of Stacey’s discretization procedure, namely as resulting from a finite elements procedure where we use a set of non-orthogonal basis functions to approximate the continuum problem. The single particle equation (1) can be obtained as the Euler-Lagrange equation corresponding to the action
$S[\psi^{\ast},\psi]$ $\!=\!\int_{t_{0}}^{t_{1}}\!\!\mathrm{d}t\int\!\!\mathrm{d}x\,\tfrac{i}{2}\psi(x,t)^{\ast}\tfrac{\partial}{\partial t}\psi(x,t)$
$\ \ -\!\tfrac{i}{2}\!\left(\tfrac{\partial}{\partial t}\psi^{\ast}(x,t)\right)\psi(x,t)\!-\!\psi(x,t)^{\ast}\!\left(\!-i\tfrac{\partial}{\partial x}\right)\psi(x,t).$
Here, the time derivative has been written in a manifestly Hermitian form, whereas the spatial derivative is Hermitian by virtue of the domain and spatial boundary conditions of $\psi$ at any instance of time.
We now express $\psi(x,t)$ with respect to a set of functions $\{\varphi_{n}(x),n\in\mathbb{Z}\}$ as
$\psi(x,t)=\sum_{n}b_{n}(t)\varphi_{n}(x)$ (7)
where the functions are not orthonormal but satisfy
$\langle\varphi_{m}|\varphi_{n}\rangle\!=\!\int\mathrm{d}x\,\varphi_{m}(x)^{\ast}\varphi_{n}(x)\!=\!N_{mn}\!=\!\begin{cases}\frac{1}{2},&n\!=\!m\\ \frac{1}{4},&n\!=\!m\!\pm\!1\\ 0,&\text{elsewhere}\end{cases}$ (8)
and
$\int\mathrm{d}x\,\varphi_{m}(x)^{\ast}\!\left(\!-i\tfrac{\partial\ }{\partial x}\right)\varphi_{n}(x)\!=\!H_{mn}\!=\!\begin{cases}+\frac{i}{2},&n\!=\!m\!+\!1\\ -\frac{i}{2},&n\!=\!m\!-\!1\\ 0,&\text{elsewhere}\end{cases}.$ (9)
An admissible set of functions that satisfy these identities in a distributional sense is given by
$\varphi_{n}(x)$ $\!=\!\tfrac{1}{\sqrt{2}}\chi_{(-1,+1)}(2x-n),$ $\!\!\!\chi_{(-1,+1)}(x)$ =$\!=\!\begin{cases}1&|x|\!<\!1\\ 0&|x|\!\geq\!1\end{cases}.$
This defines the model at the unit scale. The model at scale $\varepsilon$ can be obtained by rescaling: $\varphi^{(\varepsilon)}_{n}(x)\!=\!\tfrac{1}{\sqrt{\varepsilon}}\varphi_{n}(\varepsilon^{-1}x)$. Inserting (7) into the action gives rise to
$S[a^{\ast},a]$ $=\int_{t_{0}}^{t_{1}}\mathrm{d}t\,L(b,\dot{b},b^{\ast},\dot{b}^{\ast})$
$L$ $=\sum_{m,n}\left\{\tfrac{i}{2}\left[b_{m}^{\ast}N_{mn}\dot{b}_{n}-\dot{b}_{m}^{\ast}N_{mn}b_{n}\right]-b_{m}^{\ast}H_{mn}b_{n}\right\}.$
The canonical conjugate variables to $\{b_{n}\}$ are given by
$p_{n}=\frac{\partial L}{\partial\dot{b}_{n}}$ $=-ib_{m}^{*}N_{mn}$ (10)
We can now quantize the theory by imposing canonical anti-commutation relations
$\{p_{m},b_{n}\}=-i\delta_{m,n}$
from which we obtain [^1] In principle, all steps only hold true if $N$ is non-degenerate – otherwise, some momenta associated with the kernel of $N$ vanish. The latter situation requires the application of the Dirac algorithm for constrained systems. Alternatively, we can choose boundary conditions on finite lattices to avoid non-degeneracy of $N$, e.g., an even number of sites with anti-periodic boundary conditions or an odd number of sites with periodic boundary conditions. This does not pose any problem for the continuum and thermodynamic limit as long as 00 belongs to the continuous spectrum of $N$.
$\{a_{m},a_{n}^{\dagger}\}$ $=N_{mn},\qquad\{b_{m},b_{n}^{\dagger}\}=(N^{-1})_{mn},$
$\{a_{m},b_{n}^{\dagger}\}$ =$=\{b_{m},a_{n}^{\dagger}\}=\delta_{n,m},$
where we defined $a_{n}\!=\!b_{m}N_{mn}$ and, thus, $a_{n}^{\dagger}\!=\!b_{m}^{\dagger}N_{mn}$. The second-quantized Hamiltonian is given by
$\hat{H}$ $=\sum_{m,n}b_{m}^{\dagger}H_{mn}b_{n}=\sum_{m,n}a_{m}^{\dagger}(N^{-1}HN^{-1})_{mn}a_{n},$ (11)
to which we can now add interaction terms, as we did in Eq. (11). For the free Hamiltonian, we obtain the Heisenberg equations of motion
$\dot{a}_{k}^{\dagger}\!=\!i[\hat{H},a_{k}^{\dagger}]\!=\!i\sum_{m}b_{m}^{\dagger}H_{mk}\!=\!i\sum_{m}a_{m}^{\dagger}(N^{-1}H)_{mk},$
which entail local commutators with the Hamiltonian for observables built from $b^{{\dagger}}_{m}N_{mn}$ and $b_{m}N_{mn}$, e.g.,
$[\hat{H},b^{{\dagger}}_{m}N_{mn}b_{k}N_{kl}]$ $=b^{{\dagger}}_{m}H_{mn}b_{k}N_{kl}-b^{{\dagger}}_{m}N_{mn}b_{k}H_{kl}.$
We expect this to be relevant to prove Lieb-Robinson-type bounds in our model.
We can now build a Fock space using the creation operators $a_{m}^{\dagger}$ acting on a vacuum |$\ket{\Omega}$ with $a_{n}\ket{\Omega}=0=b_{k}\ket{\Omega}$. We expand a many-body state as
|$\ket{\Psi}=\sum_{\{n_{i}\}}C_{n_{1}n_{2}...}\,(a_{1}^{\dagger})^{n_{1}}(a_{2}^{\dagger})^{n_{2}}...\ket{\Omega}$ (12)
Denoting $C_{\vec{n}}\!=\!C_{n_{1}n_{2}...}$ and |$\ket{\vec{n}}\!=\!(a_{1}^{\dagger})^{n_{1}}(a_{2}^{\dagger})^{n_{2}}...\ket{\Omega}$, the action principle for time evolution $S[C^{\ast},C]=\int_{t_{0}}^{t_{1}}\mathrm{d}tL(t)$ with
$L(t)$ $=\sum_{\vec{n},\vec{n}^{\prime}}\left[\tfrac{i}{2}C_{\vec{n}}^{\ast}\tilde{N}_{\vec{n},\vec{n}\prime}\dot{C}_{\vec{n}^{\prime}}+h.c.\right]-C_{\vec{n}}^{\ast}\tilde{H}_{\vec{n},\vec{n}^{\prime}}C_{\vec{n}^{\prime}}$
gives rise to the equation of motion
$i\sum_{\vec{n}^{\prime}}\tilde{N}_{\vec{n},\vec{n}^{\prime}}\dot{C}_{\vec{n}^{\prime}}$ $=\sum_{\vec{n}^{\prime}}\tilde{H}_{\vec{n},\vec{n}^{\prime}}C_{\vec{n}^{\prime}}$
with $\tilde{N}_{\vec{n},\vec{n}^{\prime}}=\braket{\vec{n}}{\vec{n}^{\prime}}$ and $\tilde{H}_{\vec{n},\vec{n}^{\prime}}=\braket{\vec{n}}{\hat{H}}{\vec{n}^{\prime}}$ the retraction of the identity and Hamiltonian as introduced in the previous section. If furthermore the set of coefficients $C$, previously referred to as |$\ket{\phi}$, is also constrained to a particular form such as MPS, then we can still use the same action where we just insert this further parameterization. This gives rise to the Dirac-Frenkel time-dependent variational principle (TDVP), where now the further retraction of the norm matrix onto the tangent space of the variational manifold will appear. Importantly then, we will not have to invert the norm matrix $\tilde{N}$ in the full Hilbert space, but only its restriction onto the tangent space.
III Continuum limit
The continuum formulation of a (free) chiral fermion field describing right-moving particles can be recovered by a scaling limit procedure [23, 24]: We start with the construction in finite volume resolved at various length scales $\Lambda_{\varepsilon}\!=\!\varepsilon\{1,...,L\}\!\subset\!\varepsilon\mathbb{Z}$ ($\varepsilon L\!=\!\textup{const.}$.) as the thermodynamic limit can be achieved consecutively [25, 26]. To extract the low-energy physics of the lattice model, we focus on the many-body ground-state sector associated with (11) at a given scale $\varepsilon$, i.e., by filling up all the negative-energy modes according to the dispersion relation in (3): |$\ket{\textup{vac}_{\varepsilon}}\!\propto\!\prod_{k<0}a^{{\dagger}}_{k}\ket{\Omega_{\varepsilon}}$. The scaling limit is constructed by defining an ascending superoperator $\mathcal{R}$ identifying normal-ordered Wick monomials in $a,a^{{\dagger}}$ at different scales (following Wilson’s triangle [27, 23]). This allows us to obtain the convergence correlation functions $\bra{\textup{vac}_{\varepsilon}}a^{{\dagger}}...a...\ket{\textup{vac}_{\varepsilon}}$ at finite scales to the vacuum correlation functions $\bra{\textup{vac}_{0}}c^{{\dagger}}...c...\ket{\textup{vac}_{0}}$ of chiral fermion field $\psi$ in the continuum (with annihilation/creation operators $c,c^{{\dagger}}$). The convergence relies on the observation that the metric at scale $\varepsilon$ converges to the identity in the scaling limit, i.e., $N_{\varepsilon}\stackrel{{\scriptstyle\varepsilon\to 0}}{{\rightarrow}}\mathds{1}$. This entails that (4) reproduces the canonical anticommutation relations in the scaling limit. At the same time, the scale-dependent dispersion relation $\lambda^{(\varepsilon)}_{k}\!=\!2\varepsilon^{-1}\tan(\tfrac{1}{2}\varepsilon k)$ recovers the massless, relativistic dispersion relation $\lambda^{(0)}_{k}\!=\!k$ for $\varepsilon\to 0$. We will provide further details on the scaling limit construction elsewhere.
IV MPO representations
The (semi)definite norm matrix
$\tilde{N}_{n_{1}\ldots n_{L};n_{1}^{\prime}\ldots n_{L}^{\prime}}=\braket{\Omega}{(a_{L})^{n_{L}}\ldots(a_{1})^{n_{1}}(a_{1}^{\dagger})^{n_{1}^{\prime}}\ldots(a_{L}^{\dagger})^{n_{L}^{\prime}}}{\Omega}$
has the form of a uniform matrix product operator (MPO) with bond dimension $4$ and local tensor $\mathcal{N}$. To see this, note that a particle created at the site $i$ can be annihilated at the same and adjacent sites. This leads to six scattering types summarized in Fig. 3. Note that the type $(f)$ has an extra negative sign due to exchanging the positions of the particle. The MPO tensor $\mathcal{N}$ has the following non-zero entries (where we use the convention that the indices come in the order left, up, right, down; a 2 for the up and down index stands for a fermion present, a 1 for no fermion):
$\mathcal{N}_{1111}$ $=1,\quad\mathcal{N}_{1212}=N_{i,i}$
$\mathcal{N}_{1221}$ $=\mathcal{N}_{2112}=\mathcal{N}_{1132}=\mathcal{N}_{3211}=\sqrt{N_{i,i+1}}$
$\mathcal{N}_{1242}$ $=-\mathcal{N}_{4212}=\mathcal{N}_{2222}=\mathcal{N}_{3232}=N_{i,i+1}$
Figure 3: The $N$-dependent elements of the local norm MPO $\tilde{N}$. The vertical and horizontal bonds represent the physical and virtual spaces of dimension 2 and 4, respectively. The physical space is empty(occupied) when the index is 1(2). Fermions are able to move in a straight line or hop one site to the left or right, with weights determined by $N$. One of the entries has an extra negative sign because it exchanges the position of the two particles; the colored lines act as a visual guide.
The MPO has different boundary conditions for the even and odd particle sectors. While no boundary term is needed for the odd particle sector, in the even particle sector, we need to introduce a twist with a diagonal matrix with elements $B_{ii}=1$ for $i=1,4$ and $B_{ii}=-1$ for $i=2,3$. This can be checked by mapping $a_{i}$ to a spin operator. These boundary conditions originate from the Jordan-Wigner transformation of fermions, leading to the extra negative sign in the even particle number sector [28, 29]. Alternatively, interpreting $\mathcal{N}$ as a fermionic tensor (with physical index $2$ and virtual indices $2$ and $3$ having odd fermion parity) using the formalism from Ref. 30, no additional boundary matrix is required.
The Hamiltonian can now be constructed from this MPO by acting on it with spin creation and/or annihilation operators $\sigma^{+}=|1\rangle\langle 0|$, $\sigma^{-}=|0\rangle\langle 1|$. For example, the kinetic term which acts between sites $k$ and $k+1$ would be represented as $i\sigma^{-}_{\tilde{n}_{k},n_{k}}\tilde{N}^{n_{1}\cdots n_{L}}_{n^{\prime}_{1}\cdots n^{\prime}_{L}}\sigma^{-}_{\tilde{n}^{\prime}_{k+1}n^{\prime}_{k+1}}+h.c$
V Interacting ground state optimisation
Given the fact that both $\tilde{H}$ and $\tilde{N}$ have the form of an MPO and that the variational principle still works for generalized eigenvalue problems whenever $\tilde{N}$ is positive semidefinite (as long as the zero modes for $\tilde{N}$ are also zero-modes for $\tilde{H}$)
$\lambda_{\min}=\min_{\phi}\frac{\langle\phi|\tilde{H}|\phi\rangle}{\langle\phi|\tilde{N}|\phi\rangle}$
we can, in principle, use standard tensor network techniques to optimize the MPS [31, 32, 33]. For the non-interacting model, the filled Fermi sea results in a diverging energy density $E/L$ in the thermodynamic limit, which we expect to persist in the interacting Hamiltonian from Eq. (11). This prevents a naive implementation of the infinite lattice, but the ground state and its energy on finite periodic rings with an odd number of sites is well defined [^2] This choice of the system size prohibits the mode with momentum $\pi$. The filled Fermi sea of the non-interacting Hamiltonian also gives rise to a ground state momentum that cannot equal zero or $\pi$. Therefore, we have to use a general non-translational MPS ansatz on a ring. Similar to the standard density matrix renormalization group (DMRG) algorithm, the MPS can be optimized locally by finding the optimal tensor. More precisely, we compute an effective Hamiltonian $H_{\textup{eff}}(i)$ for site-$i$ so that $\langle\psi(A)|\tilde{H}|\psi(A)\rangle=A_{i}^{\dagger}\tilde{H}_{\textup{eff}}(i)A_{i}$. Similarly we obtain $\tilde{N}_{\textup{eff}}(i)$ as $\langle\psi(A)|\tilde{N}|\psi(A)\rangle=A_{i}^{\dagger}\tilde{N}_{\textup{eff}}(i)A_{i}$. Then, the optimal tensor $A_{i}$ can be found by solving the following linear equation
$\tilde{H}_{\textup{eff}}(i)A_{i}=\lambda_{\min}\tilde{N}_{\textup{eff}}(i)A_{i},$
where $\lambda_{\min}$ is the smallest eigenvalue corresponding to the ground state. It is important to project out the null space before inverting $\tilde{N}_{\text{eff}}$. After optimizing every site in an iterative manner, we observe that the energy converges quickly, as in conventional DRMG simulations.
Figure 4: $(a)$ The spectrum of the reduced density matrix obtained from exact diagonalization of the Hamiltonian (5) on 11 sites. The reduced density matrix is obtained by tracing out the degrees of freedom on the first five sites. $(b)$ The entanglement entropy of the free Hamiltonian ($U=0$), obtained by dividing the lattice by half on 5, 7, 9, 11, and 13 sites.
We simulate the interacting Hamiltonian (5) with $U/J=1$ and observe that the Schmidt spectrum of the halve-chain decays exponentially fast. This clearly suggests that DMRG will work equally well for the generalized eigenvalue case as for the normal case. For the case of an odd number of particles and length $L=11$; the ground state energies obtained for bond dimension $\chi=(4,6,8,10,12)$ are respectively equal to (-9.084089, -9.086027, -9.086092, -9.086096,-9.086097), which shows the fast convergence to $-9.086098$: the value obtained from the exact diagonalization [^3] We use $N_{i,i}=2,N_{i,i+1}=1$ in the following calculations. . The success of our DMRG algorithm is corroborated by the hierarchical entanglement structure of many-body ground states in one dimension. In Fig. 4, we demonstrate the spectrum of the reduced density matrix computed from exactly diagonalizing the Hamiltonian (5) on 11 sites. As is observed in conventional MPS simulations, the Schmidt values decay exponentially fast and thereby allow an efficient simulation using MPS with finite bond dimensions. $U=0$ is the gapless regime where we observe a slower decay rate. Still, the spectrum of $U>0$ can be well approximated with a few leading values. For $L=51$ and $U/J=1$, it is, of course, completely impossible to diagonalize the Hamiltonian with exact diagonalization. We obtain an energy of $-32.889236$ with a DMRG simulation with $\chi=4$.
In analogy to standard MPS techniques, developing finite-size and/or entanglement scaling methods [34, 35, 36, 37, 38, 39, 40] to extrapolate those results is significant to further improve accuracy in the future.
VI Discussion and outlook
Our results indicate that we can build lattice models of interacting chiral fermions, and efficiently target their ground state using ED [41] or MPS techniques. Developing finite-size and/or entanglement scaling methods [34, 35, 36, 37, 38, 39, 40] to extrapolate those results is significant to further improve accuracy in the future. It is also important that our method can be extended to the higher-dimensional case by using Projected Entangled Pair States (PEPS); indeed, it can readily be seen that the norm matrix in this case can efficiently be represented as a PEPO. Another question is to construct PEPS wavefunctions with the MPO $\tilde{N}$ as the leading eigenvector of the transfer matrix.
We can also generalize MPS techniques to extract the interacting excitation spectrum [42, 43] or to study dynamics. For the latter, the TDVP equation [44] seems to provide the best starting point, as the Hamiltonian $\tilde{H}$ acting on the MPS includes the norm matrix and does not admit a simple Trotter decomposition. While the effective norm matrix in the MPS tangent space (a.k.a. the metric) can no longer be made equal to the identity matrix using MPS gauge manipulations, the resulting equations can still be integrated efficiently by inverting the Gram matrix using a conjugate gradient scheme.
These algorithmic considerations are closely related to the conceptual questions regarding the locality of the generalized Schrödinger equation, which we will address in a forthcoming paper. What is the correct generalization of the Lieb-Robinson bound in this nonorthogonal quantum world, and is satisfied by a Hamiltonian that is local in the $b$ and $b^{\dagger}$ operators? Can an area law for the (Renyi) entanglement entropy in the generalized eigenvalue setting [45, 46] be proven? Another outstanding question is how to simulate our generalized Hamiltonian system on a quantum simulator or quantum computer. It is not evident to do this without coupling the system of interest to a bulk theory or by introducing long-range interactions in the form of $\tilde{N}^{-1/2}\tilde{H}\tilde{N}^{-1/2}$.
Acknowledgments: We thank Carlo Beenakker for introducing us to the concept of tangent fermions and Nick Bultinck, Karel Van Acoleyen and David Tong for helpful discussions. This work has received funding from EOS (grant No. 40007526), IBOF (grant No. IBOF23/064), BOF-GOA (grant No. BOF23/GOA/021), FWO (grant No. 12AUN24N and GOE1520N) and the MWK Lower Saxony via the Stay Inspired Program (Grant ID: 15-76251-2-Stay-9/22-16583/2022).